# s1: Simple test-time scaling

### v2

@@ Acknowledgements
This work was partly conducted using the Stanford Marlowe GPU cluster (Kapfer et al., 2025) made possible by financial support from Stanford University. We thank Alexander M. Rush, Andrew Ilyas, Banghua Zhu, Chenglei Si, Chunting Zhou, John Yang, Ludwig Schmidt, Samy Jelassi, Tengyu Ma, Xuechen Li, Yu Sun and Yue Zhang for very constructive discussions.
Appendix A Evaluation determinism
@@
Appendix B s1K details
B.1 s1K summary
@@
B.2 Dataset composition for full 59K questions
@@
B.3 s1K grading prompt
To grade whether an example is correct for our dataset selection in §2, we use the prompt in Figure 8.
@@ B.5 Decontamination
Appendix C Training details
We take a model that has already been pretrained and instruction tuned and further finetune it for reasoning. Specifically, we use Qwen2.5-32B-Instruct (Qwen et al., 2024), which on math tasks generally matches or outperforms the larger Qwen2.5-72B-Instruct (Qwen et al., 2024) or other open models (Dubey et al., 2024; Groeneveld et al., 2024; Muennighoff et al., 2024). We use token delimiters to separate the thinking stage from the answering stage. We enclose the thinking stage with <|im_start|>think and <|im_start|>answer; both preceded and followed by a newline. Samples from our dataset are in §C.2. We use basic fine-tuning hyperparameters: we train for 5 epochs with a batch size of 16 for a total of 315 gradient steps. We train in bfloat16 precision with a learning rate of $51e-$ 5 warmed up linearly for 5% (16 steps) and then decayed to 0 over the rest of training (299 steps) following a cosine schedule. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9,\beta_{2}=0.95$ and weight decay of $41e-$ 4. We do not compute loss on questions, only on reasoning traces and solutions. We ensure the sequence length is large enough to avoid cutting off any samples; a setting we ablate in §C.1. The training takes just 26 minutes on 16 NVIDIA H100 GPUs.
@@ Appendix C Training details
C.1 Training Ablations: Sequence length
@@ C.2 Training Samples
Appendix D Test-time scaling details
D.1 Sequential scaling ablations
@@
D.2 Examples for rejection sampling ablation

### v3

@@ Acknowledgements
We thank Ryan Marten for generating traces from DeepSeek r1 for s1.1 using Bespoke Curator (Marten et al., 2025). This work was partly conducted using the Stanford Marlowe GPU cluster (Kapfer et al., 2025), made possible by financial support from Stanford University. We thank Alexander M. Rush, Andrew Ilyas, Banghua Zhu, Chenglei Si, Chunting Zhou, John Yang, Ludwig Schmidt, Samy Jelassi, Suhas Kotha, Tengyu Ma, Xuechen Li, Yu Sun, and Yue Zhang for very constructive discussions.
Appendix A s1.1
Seven days after our release of s1, we released s1.1. We regenerated traces for our 1,000 samples in s1K using DeepSeek r1 (DeepSeek-AI et al., 2025) to create s1K-1.1. We use the same training procedure to train our model s1.1. Other updates since our launch include the release of o3 (OpenAI, 2025), LIMO (Ye et al., 2025a), and AIME 2025. We consider all these new developments in LABEL:tab:perf2. We find that s1.1 performs significantly better than s1. We also tried distilling from Claude 3.7, which led to worse performance than from r1 (not reported).
Table 5: s1-32B is an open and sample-efficient reasoning model. We evaluate s1-32B, Qwen, and Gemini (some entries are unknown (N.A.), see §4). Other results are from the respective reports (Qwen et al., 2024; Team, 2024; OpenAI, 2024, 2025; DeepSeek-AI et al., 2025; Labs, 2025; Team, 2025) except for AIME 2025 (Ye et al., 2025b). # ex. = number examples used for reasoning finetuning; BF = budget forcing.
Model # Examples MATH500 GPQA AIME 2024 AIME 2025
API only
o3-mini-low N/A 95.8 70.6 56.3 42.1
o3-mini-medium N/A 97.3 76.8 75.8 70.4
o3-mini-high N/A 97.9 79.7 83.8 80.9
Open Weights
QwQ-32B N.A. 90.6 54.5 46.7 32.7
r1 $\gg$800K 07.3 71.5 79.8 70.0
r1-distill-Llama-70B 800K 94.5 65.2 57.1 56.3
r1-distill-Qwen-14B 800K 93.9 59.1 61.7 48.0
r1-distill-Qwen-32B 800K 94.3 62.1 58.3 49.6
Open Weights and Open Data
LIMO 817 94.8 66.7 56.3 44.6
s1 w/o BF 1K 92.6 56.6 50.0 26.7
s1 with Budget Forcing “Wait” 1x 1K 92.8 59.6 53.3 30.0
s1 with Budget Forcing “Wait” 2x 1K 93.0 59.6 53.3 33.3
s1 with Budget Forcing “Wait” 4x 1K 92.2 58.6 56.7 36.7
s1.1 w/o BF 1K 94.4 60.6 56.7 50.0
s1.1 with Budget Forcing “Wait” 1x 1K 95.4 62.6 56.7 50.0
s1.1 with Budget Forcing “Wait” 2x 1K 95.4 63.6 56.7 50.0
Appendix B Evaluation determinism
@@
Appendix C s1K details
C.1 s1K summary
@@
C.2 Dataset composition for full 59K questions
@@
C.3 s1K grading prompt
To grade whether an example is correct for our dataset selection in §2, we use the prompt in Figure 8. We grade using Claude 3.5 except for the correctness among the final 1,000 samples, which we graded with Claude 3.7.
@@ C.5 Decontamination
Appendix D Training details
We take a model that has already been pretrained and instruction tuned and further finetune it for reasoning. Specifically, we use Qwen2.5-32B-Instruct (Qwen et al., 2024), which on math tasks generally matches or outperforms the larger Qwen2.5-72B-Instruct (Qwen et al., 2024) or other open models (Dubey et al., 2024; Groeneveld et al., 2024; Muennighoff et al., 2024). We use token delimiters to separate the thinking stage from the answering stage. We enclose the thinking stage with <|im_start|>think and <|im_start|>answer; both preceded and followed by a newline. Samples from our dataset are in §D.2. We use basic fine-tuning hyperparameters: we train for 5 epochs with a batch size of 16 for a total of 315 gradient steps. We train in bfloat16 precision with a learning rate of $51e-$ 5 warmed up linearly for 5% (16 steps) and then decayed to 0 over the rest of training (299 steps) following a cosine schedule. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9,\beta_{2}=0.95$ and weight decay of $41e-$ 4. We do not compute loss on questions, only on reasoning traces and solutions. We ensure the sequence length is large enough to avoid cutting off any samples; a setting we ablate in §D.1. The training takes just 26 minutes on 16 NVIDIA H100 GPUs.
@@ Appendix D Training details
D.1 Training Ablations: Sequence length
@@ D.2 Training Samples
Appendix E Test-time scaling details
E.1 Sequential scaling ablations
@@
E.2 Examples for rejection sampling ablation
@@
Appendix F Version Control
V2 → V3 (2025-03):
  • Added §A
  • Added number of correct samples in §2
V1 → V2 (2025-02):
  • Added citations and other small writing changes